Write v = (2, -5 , 3) as a linear combination of

u1=(1,−3 ,2)

u2=(2,−4 ,−1)

u3=(1,−3 , 7)


1
Expert's answer
2021-11-04T20:15:34-0400

Let "x_1u_1+x_2u_2+x_3u_3=v\quad x\in\reals"


"\begin{bmatrix}
 1&2&1 \\
 -3&-4&-3\\2&-1&7
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2\\x_3
\end{bmatrix}=\begin{bmatrix}
 2 \\-5\\3
 
\end{bmatrix}"


Considering augmented matrix "A" for this system and applying Gauss- Jordan elimination


"A=\begin{matrix}
 |&1&2&1\\|&-3&-4&-3&\\|&2&-1&7
\end{matrix}\begin{vmatrix}
 2\\-5\\3
 
\end{vmatrix}"


"-\frac{1}{3}" "R_2-R_1\to\>R_2"


"\begin{matrix}
 |&1&2&1 \\
 |&0&-\frac{2}{3} & 0\\
|&1&-\frac{1}{2}&\frac{7}{2}
\end{matrix}\begin{vmatrix}
 2\\
 -\frac{1}{3}\\
\frac{3}{2}
\end{vmatrix}"



"\frac{1}{2}\>R_3-R_1\>\to\>R_2"



"\begin{matrix}
 |&1&2&1 \\
 |&0&-\frac{2}{3}&0 \\
|&0&-\frac{5}{2}&\frac{5}{2}
\end{matrix}\begin{vmatrix}
 2 \\
 -\frac{1}{3} \\
-\frac{1}{2}
\end{vmatrix}"




"-\frac{3}{2}\>R_2\to\>R_2" "\begin{matrix}
 |&1&2&1 \\
 |&0&1&0 \\
|&0&-\frac{5}{2}&\frac{5}{2}
\end{matrix}\begin{vmatrix}
 2 \\
 \frac{1}{2} \\
-\frac{1}{2}
\end{vmatrix}"


"-\frac{2}{5}R_3-R_2\to\>R_3"


"\begin{matrix}
 |&1&2&1 \\
 |&0&1&0 \\
|&0&0&-1
\end{matrix}\begin{vmatrix}
 2 \\
 \frac{1}{2} \\
-\frac{3}{10}
\end{vmatrix}"



"-1" "R_3\to\>R_3" "\begin{matrix}
 |&1&2&1 \\
 |&0&1&0 \\
|&0&0&1
\end{matrix}\begin{vmatrix}
 2 \\
 \frac{1}{2}\\
\frac{3}{10}
\end{vmatrix}"



"R_1-R_3\to\>R_1" "\begin{matrix}
 |&1&2&0 \\
 |&0&1&0\\
|&0&0&1
\end{matrix}\begin{vmatrix}
 2 \\
 \frac{1}{2} \\
\frac{3}{10}
\end{vmatrix}"



"R_1-2R_2\to\>R_1" "\begin{matrix}
 |&1&0&0 \\
 |&0&1&0 \\
|&0&0&1
\end{matrix}\begin{vmatrix}
 \frac{7}{10} \\
 \frac{1}{2} \\
\frac{3}{10}
\end{vmatrix}"


rref "A=" "\begin{matrix}|&
 1&0&0 \\|&
 0&1&0\\|&0&0&1
\end{matrix}\begin{vmatrix}
 \frac{7}{10}\\\frac{1}{2}\\\frac{3}{10}
 
\end{vmatrix}"


"\implies x_1=\frac{7}{10}\quad x_2=\frac{1}{2} \quad x_3=\frac{3}{10}"



"v=\frac{7}{10}u_1+\frac{1}{2}u_2+\frac{3}{10}u_3"